# On the Compositional Generalization of Multimodal LLMs for Medical Imaging

### v1

@@ 3.2.1 Experiment Setup
3.2.2 Analysis of Scaling Experiment
@@
Besides, w/o Modality / Area / Task showed significant accuracy drops compared to All Related, despite holding the training data volume constant. Notably, All Related performs at a level comparable to All Data, where all datasets are included in training. This suggests that CG provides significant help for the generalization brought by multi-task training.
@@
Take-away 5: CG is one of the primary forms of generalization for MLLMs trained on multi-task dataset in medical imaging (Q2).
| Related Combination | Target Subset | Baseline | Trained | ||
| CT - Subset02 | Brain - Subset22 | Cancer - Subset07 | CT, Brain, Cancer | 28 | 26 |
| CT - Subset03 | Brain - Subset22 | Cancer - Subset21 | CT, Brain, Cancer | 28 | 25 |
| CT - Subset02 | Brain - Subset22 | State - Subset09 | CT, Brain, State | 33 | 64 |
| CT - Subset03 | Brain - Subset22 | State - Subset26 | CT, Brain, State | 33 | 70 |
| X-ray - Subset25 | Lung - Subset03 | Diseases - Subset02 | X-ray, Lung, Diseases | 30 | 45 |
| X-ray - Subset26 | Lung - Subset03 | Diseases - Subset02 | X-ray, Lung, Diseases | 30 | 38 |
| X-ray - Subset26 | Lung - Subset03 | Diseases - Subset08 | X-ray, Lung, Diseases | 30 | 44 |
| X-ray - Subset26 | Breast - Subset24 | Diseases - Subset02 | X-ray, Breast, Diseases | 31 | 32 |
| X-ray - Subset28 | Breast - Subset24 | Diseases - Subset08 | X-ray, Breast, Diseases | 31 | 52 |
4 In-depth Analysis of CG
To further explore the application and applicability of CG, several research questions were proposed in this section. In order to present rich results with minimal consumption of computational resources, specific data Selection Strategies were applied for each question. Here are the research questions:
(RQ 1): How does the quantity of Related data affect generalization, and should it be maximized to ensure CG quality?
(RQ 2): Can CG combinations help the model efficiently fit Target data when only a small amount of Target data is available?
(RQ 3): Previous studies demonstrated that CG exists between two MAT-Triplet elements, does CG still exist if the three MAT-Triplet elements come from three different datasets?
(RQ 4): Does CG exist between images from Detection tasks and Classification tasks?
(RQ 5): Can different MLLM backbones leverage CG to interpret unseen medical images?
4.1 The Relationship Between CG Performance and Data Volume
@@
Selection Strategy: This section focuses on examining how the data volume of Related combinations influences generalization. To highlight the generalization trends, the combinations with strong generalization results were selected from the main experiments. For fairness, we chose the combinations across four types where Trained results exceed both Baseline and Baseline+ by at least 10. If multiple combinations meet the criteria, a random seed of 42 was used to determine the selection.
In this experiment, the amount of combination data was gradually increased (from 0 to 750, 1500, 2250, and 3000), and the trend of the model’s accuracy on Target data after training was observed. The experimental results are shown in Figure 5, where the red line represents the accuracy curve gained from Related combinations, and the purple line represents the gain from Unrelated combinations. Models trained on Related combinations demonstrate strong performance across all four Target data sets. As the data volume of Related combinations increases, the model’s understanding of Target data improves consistently.
Take-away 6: Generalization effect improves with increasing CG combination’s volume. (RQ 1)
4.2 Achieving Data-Efficient Training through CG
Selection Strategy: This experiment evaluates how CG aids training and dataset fitting with limited Target data. To highlight performance trends, four combinations with poor generalization (marked in red in Table 3) were selected across the four Direction Types, ensuring both models in the control group started with similar baseline accuracies. When multiple poor-performing combinations existed within a Direction Type, the more challenging datasets were chosen based on Table 1 to better observe potential improvements.
In this experiment, the amount of Target data was progressively increased (from 0 to 500, 1000, 1500, and 2000), and a fixed number of combination data were incorporated during training. Figure 6 displays the results, with red lines showing accuracy gains from Related combination and purple lines from Unrelated ones. It can be observed that training with Related combination data reaches peak performance more quickly, demonstrating another form of CG. Therefore, although CG performs poorly on these data due to task characteristics, it still helps the model quickly adapt with only a small amount of Target data.
@@ 4.2 Achieving Data-Efficient Training through CG
Take-away 7: Although CG does not always provide direct generalization gains, it can help Target data achieve data-efficient training. (RQ 2)
4.3 CG with All MAT-Triplet Elements from Different Sources
In previous controlled experiments (Section 3.1), one element of the MAT-Triplet was kept constant while CG was explored in the remaining two elements. To ensure that all the 3 MAT-Triplet elements of the target data originated from three distinct datasets, additional experiments were conducted to further validate the effectiveness of CG. For these experiments, all possible combinations meeting the criteria in Med-MAT were selected (Selection Strategy). The results presented in Table 4 demonstrate that most combinations can effectively generalize to the Target data.
Take-away 8: CG remains effective when all elements come from different datasets (RQ 3).
4.4 Detection-Capable MLLMs can Leverage Detection Data for CG
@@
Previous studies Ren et al. (2024); Wang et al. (2025) have demonstrated that jointly training classification and detection tasks can mutually enhance their performance. Building on this, we explored CG within this context. To ensure comprehensive conclusions, we conducted experiments using Next-Chat and MiniGPT-v2, which represent the two mainstream approaches for MLLMs to perform segmentation tasks. The former treats bounding boxes as embeddings and decodes them into coordinates using an additional visual decoder, while the latter processes coordinate points as special text tokens and generates bounding box coordinates directly as output text. The final results (Figure 7) show that all CG combinations demonstrated the model’s successful utilization of detection data for CG to the Target data.
@@
Take-away 9: MLLMs can combine knowledge from detection tasks with classification tasks to improve classification accuracy (RQ 4).
4.5 CG Exists Across Different MLLM Backbones
LLaVA was selected as the baseline because its training data and processes are publicly available, ensuring minimal exposure to medical images and preventing bias in the integration of medical image knowledge into the MLLM. To verify that CG is not limited to a single framework, we also selected subsets for experiments on Qwen2-VL-7B Wang et al. (2024a) and Llama3.2-11B-Vision Meta AI (2024). The results in Figure 7 demonstrate CG effects similar to those observed in the LLaVA.
Take-away 10: CG persists across different MLLM backbones (RQ 5).
@@ 5 Related Work
Medical MLLMs: Recently, adapting MLLMs to medical tasks has gained prominence due to their success in capturing complex visual features. Current MLLMs typically pair a visual encoder with a text-only LLM, aligning image data with language understanding. Such as Med-Flamingo (Moor et al., 2023) and Med-PaLM (Tu et al., 2024), fine-tuned general multimodal models and achieved notable results. Med-Flamingo enhanced OpenFlamingo-9B (Chen et al., 2024a) with medical data, while Med-PaLM adapted PaLM-E (Driess et al., 2023) using 1 million data points. Similarly, LLaVA-Med Li et al. (2024), Med-Gemini (Saab et al., 2024), and HuatuoGPT-Vision Chen et al. (2024b) utilized specialized datasets and instruction tuning to refine medical VQA tasks.
Generalization on Medical Imaging: Generalization in medical imaging Matta et al. (2024) has been extensively studied. Early methods utilized data manipulation techniques, such as data augmentation Li et al. (2022); Zhang et al. (2022), to enhance model generalization on unseen medical data by adapting to varying distributions. Later approaches focused on representation learning Le-Khac et al. (2020), preserving essential image information to enable models to handle more complex scenarios. Additionally, some studies Ren et al. (2024) explore multiple aspects of medical image processing, examining how classification and segmentation tasks can mutually benefit each other.
Detection with MLLMs: MiniGPT-4 Zhu et al. (2023) and LLaVA Liu et al. (2023) enhance LLMs by fine-tuning models on synthetic multimodal instruction data. However, these models are limited to image and text inputs and cannot generate detection bounding boxes or segmentation masks. Recent works have addressed these limitations using various strategies, such as encoding regions as features to allow models to accept regions as input Zhang et al. (2023b), representing object bounding box coordinates with text tokens Wang et al. (2024c); Peng et al. (2023); Chen et al. (2023b), and employing unique identifiers for task instructions to improve learning efficiency. Additionally, some approaches introduce special tokens to represent images and use their hidden states to decode position information Zhang et al. (2023a, 2024a).
@@ 6 Conclusion
To explore whether MLLMs can leverage CG to generalize unseen medical data, we constructed the Med-MAT dataset as the research platform for generalization experiments. The results confirmed the existence of CG and identified it as one of the primary factors behind the generalization capability of MLLMs in multi-task learning. Subsequent experiments demonstrated that increasing the volume of CG combinations consistently improved their effectiveness. Furthermore, CG enables medical tasks to efficiently fit with limited data, reducing dependence on the volume of training data. Importantly, CG is present across various MLLM backbones and can even facilitate generalization using data from tasks such as detection, underscoring its broad applicability.
@@
The experiment confirms that MLLMs can use CG to understand unseen medical images and achieve data-efficient training. However, as shown in the results in 3.2, even after disrupting CG, the model’s generalization performance declines but still maintains some level of effectiveness. Thus, CG represents only one form of generalization for MLLMs in medical imaging. To support the exploration of other generalization mechanisms, we will make datasets publicly available according to the license.
This study is focused exclusively on medical scenarios, but we believe that similar generalization effects may also apply in other multimodal tasks. Additionally, more granular medical generalization strategies, such as leveraging existing pneumonia data to generalize the detection of newly emerging types of pneumonia, are also promising areas for further investigation.
@@ Appendix A More Experiments
A.1 CG on Detection Tasks (RQ 4)
| Related Combination | Target Subset | Baseline | Trained | |
| Lung, Lung Det | Bones, Diseases | Lung, Diseases | 49 | 52 |
| Lung, Lung Det | Breast, Diseases | Lung, Diseases | 49 | 54 |
| Bones, Spinal Error Det | Breast, Diseases | Bones, Diseases | 20 | 30 |
| Bones, Spinal Error Det | Lung, Diseases | Bones, Diseases | 20 | 33 |
| MRI, Diseases Det | End, Level | End, Diseases | 24 | 27 |
| X-ray, Lung Det | CT, COVID | X-ray, COVID | 23 | 26 |
| Der, Skin, Cancer Det | FP, Fundus, Diseases | Der, Skin, Diseases | 24 | 29 |
| Mic, Cell, Cancer Det | CT, Kidney, Diseases | Mic, Cell, Diseases | 24 | 26 |
| Related Combination | Target Subset | Baseline | Trained | |
| Lung, Lung Det | Bones, Diseases | Lung, Diseases | 41 | 47 |
| Lung, Lung Det | Breast, Diseases | Lung, Diseases | 41 | 49 |
| Bones, Spinal Error Det | Breast, Diseases | Bones, Diseases | 31 | 35 |
| Bones, Spinal Error Det | Lung, Diseases | Bones, Diseases | 31 | 37 |
| MRI, Diseases Det | End, Level | End, Diseases | 24 | 26 |
| X-ray, Lung Det | CT, COVID | X-ray, COVID | 22 | 23 |
| Der, Skin, Cancer Det | FP, Fundus, Diseases | Der, Skin, Diseases | 27 | 30 |
| Mic, Cell, Cancer Det | CT, Kidney, Diseases | Mic, Cell, Diseases | 20 | 24 |
Previous studies have shown that training on detection datasets can improve model performance in classification tasks. Building on this insight, we further examine its applicability to CG. For this investigation, all possible detection subset combinations in Med-MAT were selected to explore their impact on CG (Selection Strategy).
The two backbone was trained separately in this experiment. For Next-Chat, we directly trained the model in its second training stage and fine-tuned it for 2 epochs on our composition datasets with a learning rate of 2e-5, keeping all other training parameters at their default settings. Similarly, for MiniGPT-v2, we trained the backbone model from the second stage, starting with a learning rate of 2e-5 and gradually reducing it to 2e-6 over 3 epochs.
The experiments in Table 5 and Table 6 present the results for Next-Chat and MiniGPT-v2, respectively. The separation regions closely align with the 4 Direction Types defined in the classification task, except for Fix Task, which is excluded due to the substantial differences between detection and classification tasks, making it impossible to find the fix one. The results demonstrate that joint training on detection and classification tasks effectively compositional generalizes to classification tasks in both backbones, leading to higher model accuracy. This indicates that CG is also present in these tasks.
| Related Combination | Target Subset | Baseline | Trained | |
| Bones, State | Breast, Diseases | Bones, Diseases | 61 | 65 |
| Lung, COVID | Bones, Diseases | Lung, Diseases | 80 | 91 |
| CT, COVID | X-ray, Diseases | X-ray, COVID | 35 | 40 |
| CT, State | X-ray, Diseases | X-ray, State | 35 | 43 |
| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 32 | 33 |
| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 65 | 72 |
| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 48 | 45 |
| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 34 | 41 |
| Related Combination | Target Subset | Baseline | Trained | |
| Bones, State | Breast, Diseases | Bones, Diseases | 52 | 59 |
| Lung, COVID | Bones, Diseases | Lung, Diseases | 64 | 75 |
| CT, COVID | X-ray, Diseases | X-ray, COVID | 33 | 38 |
| CT, State | X-ray, Diseases | X-ray, State | 33 | 41 |
| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 31 | 29 |
| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 49 | 57 |
| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 55 | 61 |
| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 10 | 32 |
A.2 Other MLLMs (RQ 5)
The choice to use models like LLaVA, MiniGPT-v2 and NExT-GPT is because their training data and overall process are publicly available, ensuring that they have not been too much exposed to medical images, thus preventing bias in the medical image knowledge injection into the MLLM. To ensure the experiment results are not influenced by the model choice, we also tested several other models on some subsets of Med-MAT and observed similar results.

### v2

@@ 4.1 Experiment Setup
4.2 Analysis of Scaling Experiment
@@
Besides, w/o Modality / Area / Task showed significant accuracy drops compared to All Related, despite holding the training data volume constant. This indicates that if the CG combinations are forcibly disrupted, MLLMs will lose a significant amount of generalization capability for the target data.
Take-away 5: Disrupting CG leads to a significant decline in generalization ability. (Q2).
Notably, All Related achieves a performance level comparable to All Data, where all datasets are included in training. This suggests that CG plays a crucial role in enhancing the generalization effect of Multi-task Training. Therefore, in conclusion:
@@
Take-away 6: CG plays an important role in generalization for MLLMs in medical imaging.
5 Potential Applications of CG
As MLLMs can use CG to generalize unseen medical images, this section attempts to explore its potential applications in training medical MLLMs.
5.1 Generalization without Target Data
In medical tasks, new and unpredictable conditions, like COVID-19, can emerge at any time. Exploring how to use CG to help MLLMs enhance their ability to identify unknown diseases in the absence of specific datasets is both important and meaningful.
We selected some Target datasets and trained the MLLMs using Related and Unrelated data to observe their generalization to the Target data. The generalization trend was assessed by progressively increasing the size of the combination datasets.
Selection Strategy
@@ 5.1 Generalization without Target Data
To highlight the generalization trends, the combinations with strong generalization results were selected from the main experiments. For fairness, we chose the combinations across four types where Trained results exceed both Baseline and Baseline+ by at least 10. If multiple combinations meet the criteria, a random seed of 42 was used to determine the selection.
Analysis
The experimental results are shown in Figure 5, where the red line represents the accuracy curve for Related combinations, and the purple line shows the gain from Unrelated combinations. The Related combinations group significantly outperformed the Unrelated combinations in terms of generalization across all tasks, with this ability continuing to improve as the data size increased. This suggests that Related combinations, leveraging CG, enhance the model’s ability to understand unknown medical tasks.
Take-away 7: CG might enable MLLMs to handle tasks without dedicated training data.
5.2 Generalization with Limited Target Data
This section investigates the benefit of CG for tasks with limited data, e.g. processing medical images in rare conditions.
Selection Strategy
To assess generalization in limited data scenarios, we select combinations with poor generalization from Table 3. Specifically, for each Direction Type, we randomly choose a CG combination with weak generalization (i.e., rows marked with ✗ in the last column of Table 3). For these combinations, we introduce an additional 2,000 examples from the Target data.
Analysis
Figure 6 shows the results. It can be seen that as we gradually expand the training volume of Target data, adding the Related combination for training enabled the model to reach the peak performance more quickly. This suggests that leveraging CG to assist low-data medical scenarios can lead to more data-efficient training, even when CG does not directly result in significant generalization gains in these scenarios.
@@ 5.2 Generalization with Limited Target Data
Take-away 8: Although CG might not provide direct generalization gains, it helps data efficiency for MLLM training.
6 CG across Detection and Classification
Previous studies Ren et al. (2024); Wang et al. (2025) have shown that jointly training classification and detection tasks can mutually enhance their performance. Building on this, we investigate whether MLLMs can leverage classification data (e.g., visual knowledge) and detection data (e.g., spatial information) through CG to improve downstream classification (Q1) or detection tasks (Q2).
6.1 Experiment Settings
Training Setup
Each generalization combination used for training in this experiment includes one detection dataset and one classification dataset to examine the generalization relationship between these two vision tasks. The detailed training parameters can be found in Appendix A.6.
Model Selection
@@ 6.1 Experiment Settings
Next-Chat Zhang et al. (2023a) and MiniGPT-v2 Chen et al. (2023a) are selected as baselines, representing the two main approaches MLLMs use for detection tasks. The former treats bounding boxes as embeddings and decodes them into coordinates using a visual decoder, while the latter processes coordinate points as special text tokens and generates bounding box coordinates directly as output text.
Data Processing
Med-MAT includes both detection and segmentation datasets. If a segmentation dataset provides object localization using masks, we extract the outermost coordinates of the corresponding mask to construct a bounding box, facilitating generalization experiments for detection. Subsequently, to streamline the experiments, we structured the dataset following the official data formats of Next-Chat and MiniGPT-v2.
6.2 Benefits for Classification (Q1)
In this experiment, all possible CG combinations were selected and the CG-trained model will be tested on classification task. The final results in Figure 7 show that all CG combinations demonstrated the model’s successful utilization of detection data for CG to the Target data.
6.3 Benefits for Detection (Q2)
Subset 38 and 39 are selected as the objects in these datasets are relatively randomly distributed in the images, making them suitable for evaluating the model’s detection capability. Subsequently, we selected certain classification datasets to construct CG for testing and used cIoU to evaluate the detection performance (follow Chen et al. (2023a)).
Since both baselines lack localization capabilities for medical tasks, we incorporated a fixed amount of Target data into our experiments, adjusting the evaluation scenario to assess support in low-data settings. The results in Table 5 show that all selected CG combinations help MLLMs achieve better performance in detection tasks.
| Related Combination | Target Subset | Next-Chat | MiniGPT-v2 | |
| - Skin | - Intestine | - Intestine | ||
| - Intestine | - Skin | - Skin | ||
@@ 6.3 Benefits for Detection (Q2)
Take-away 9: MLLMs can perform CG across classification and detection tasks.
@@ 7 Related Work
Medical MLLMs Recently, adapting MLLMs to medical tasks has gained prominence due to their success in capturing complex visual features. Current MLLMs typically pair a visual encoder with a text-only LLM, aligning image data with language understanding. Such as Med-Flamingo (Moor et al., 2023) and Med-PaLM (Tu et al., 2024), fine-tuned general multimodal models and achieved notable results. Med-Flamingo enhanced OpenFlamingo-9B (Chen et al., 2024a) with medical data, while Med-PaLM adapted PaLM-E (Driess et al., 2023) using 1 million data points. Similarly, LLaVA-Med Li et al. (2024), Med-Gemini (Saab et al., 2024), and HuatuoGPT-Vision Chen et al. (2024b) utilized specialized datasets and instruction tuning to refine medical VQA tasks.
Generalization on Medical Imaging
Generalization in medical imaging Matta et al. (2024) has been extensively studied. Early methods utilized data manipulation techniques, such as data augmentation Li et al. (2022); Zhang et al. (2022), to enhance model generalization on unseen medical data by adapting to varying distributions. Later approaches focused on representation learning Le-Khac et al. (2020), preserving essential image information to enable models to handle more complex scenarios. Additionally, some studies Ren et al. (2024) explore multiple aspects of medical image processing, examining how classification and segmentation tasks can mutually benefit each other.
Detection with MLLMs
Recent studies employ various strategies to equip MLLMs with the capability to handle detection tasks, such as encoding regions as features to allow models to accept regions as input Zhang et al. (2023b), representing object bounding box coordinates with text tokens Wang et al. (2024c); Peng et al. (2023); Chen et al. (2023b), and employing unique identifiers for task instructions to improve learning efficiency. Additionally, some approaches introduce special tokens to represent images and use their hidden states to decode position information Zhang et al. (2023a, 2024a).
@@ 8 Conclusion
To investigate whether MLLMs can leverage CG to generalize to unseen medical data, we constructed the Med-MAT dataset as a research platform for generalization experiments. The results confirmed the presence of CG and identified it as a key factor of MLLMs’ generalization observed in multi-task learning. Further experiments showed that CG helps MLLMs handle limited data conditions, providing support for low-data medical tasks. Additionally, our findings showed that MLLMs can apply CG across detection and classification tasks, underscoring its broad generalization potential.
@@
The experiment confirms that MLLMs leverage CG for unseen medical images and data-efficient training. However, as shown in Section 4, disrupting CG reduces generalization but retains some effectiveness, indicating CG is just one aspect of MLLM generalization in medical imaging.
@@
Acknowledgments
This work was supported by Shenzhen Medical Research Fund (No.C2406002) from the Shenzhen Medical Academy of Research and Translation (SMART), the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Science and Technology Program (Shenzhen Key Laboratory Grant No. ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
@@ Appendix A More Experiments
A.1 Benefits for Segmentation
Segmentation-enabled LLMs, such as Next-GPT, first use the LLM to identify potential regions of the target object and then apply a SAM to decode the object mask, thereby completing the segmentation task. In this context, segmentation can be seen as an extension of detection, potentially requiring more images to achieve improved performance. We conducted additional experiments to explore whether MLLMs can still utilize CG to understand new images across both segmentation and classification tasks.
| Related Combination | Target Subset | Next-Chat | |
| - Skin | - Intestine | - Intestine | |
| - Intestine | - Skin | - Skin | |
The results in Table 6 demonstrate that, in the context of segmentation tasks, MLLMs are still able to leverage CG to understand new tasks, which is consistent with our original conclusions.
A.2 More Complex Medical Elements
While MAT-Triplet Categorization is useful, predefined categories may limit the exploration of more complex medical attributes, so we also considered integrating more flexible categorization to explore additional medical attributes.
Additional Element 1: Population Groups
We selected VinDr-PCXR and MedMAT Subset 31 for the experiment, as they contain X-ray images of children and adult groups, respectively. The results are shown in Table 7.
Additional Element 2: Finer Disease
"Finer disease" means more detailed categorization. For instance, we treat COVID and common pneumonia as distinct diseases for generalization. We split the Normal data in the training set into two parts and combined each with COVID and Pneumonia data to create new datasets. The results are shown in Table 8.
| Related Combination | Target Subset | LLaVA | ||||
Children (CG) |
| Related Combination | Target Subset | LLaVA | ||||
| Related Combination | Target Subset | Baseline | 1st | 2nd | 3rd | Mean and SD | ||||
A.3 Statistical Tests of the Generalization Results
To ensure consistency and repeatability of the experiment, we performed statistical tests in this section. LLaVA is selected as the baseline, and we used the same data combinations from Section 3.3. Each experiment was repeated 3 times, and we reported the mean and standard deviation (SD) of the results.
From the results in Table 9, we can observe that the outcomes across runs show low variance, indicating overall stability, and they continue to support our original experimental conclusions.
A.4 CG with All MAT-Triplet Elements from Different Sources
In previous controlled experiments (Section 3), one element of the MAT-Triplet was kept constant while CG was explored in the remaining two elements. To ensure that all the 3 MAT-Triplet elements of the target data originated from three distinct datasets, additional experiments were conducted to further validate the effectiveness of CG. For these experiments, all possible combinations meeting the criteria in Med-MAT were selected (Selection Strategy). The results presented in Table 10 demonstrate that most combinations can effectively generalize to the Target data.
| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
28 | 26 | ✗ |
28 | 25 | ✗ |
33 | 64 | ✓ |
33 | 70 | ✓ |
30 | 45 | ✓ |
30 | 38 | ✓ |
30 | 44 | ✓ |
31 | 32 | ✓ |
31 | 52 | ✓ |
| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
61 | 65 | ✓ |
80 | 91 | ✓ |
35 | 40 | ✓ |
35 | 43 | ✓ |
32 | 33 | ✓ |
65 | 72 | ✓ |
48 | 45 | ✗ |
34 | 41 | ✓ |
| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
52 | 59 | ✓ |
64 | 75 | ✓ |
33 | 38 | ✓ |
33 | 41 | ✓ |
31 | 29 | ✗ |
49 | 57 | ✓ |
55 | 61 | ✓ |
10 | 32 | ✓ |
Analysis of the results
The results in Table 7 and 8 indicate that the two new attributes show data leakage due to subtle visual differences in corresponding images (e.g., COVID-19 and pneumonia have similar features). Importantly, the MLLM trained with CG combinations still shows improvements on downstream tasks, confirming that our approach remains valid for new attributes.
Reason to choose the existing three attributes (MAT-Triplet: Modality, Area, Task)
We have considered additional categories such as age, gender, and finer disease classification, but we ultimately chose to focus on the MAT-Triplet categories for the following reasons.
The boundaries between MAT-Triplet (Modality, Area, Task) are clear. Different modalities and areas correspond to distinct imaging methods and body areas, leading to significant differences between images; different tasks also require the MLLM to extract specific information, demanding varied understanding of the images.
All datasets can be annotated using MAT-Triplet (Modality, Area, Task) easily. Other medical labels, such as gender and age, are only available in a small portion of datasets and are not suitable for large-scale annotation.
Similar categorization strategies have been adopted in previous studies.
| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
49 | 52 | ✓ |
49 | 54 | ✓ |
20 | 30 | ✓ |
20 | 33 | ✓ |
24 | 27 | ✓ |
23 | 26 | ✓ |
24 | 29 | ✓ |
24 | 26 | ✓ |
| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
41 | 47 | ✓ |
41 | 49 | ✓ |
31 | 35 | ✓ |
31 | 37 | ✓ |
24 | 26 | ✓ |
22 | 23 | ✓ |
27 | 30 | ✓ |
20 | 24 | ✓ |
A.5 Details of Section 3.3: Exploring CG on different MLLM Backbones
To ensure the experiment results are not influenced by the model choice, we also tested several other models on some subsets of Med-MAT and observed similar results.